OpenStudy (anonymous):

need calculus question check. 1 sec, let me rewrite question.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} \sin^{-1} (x ^{2}+x-2)/6(x-1) \rightarrow \lim_{x \rightarrow 1} \sin^{-1} ((x-1)(x-1))/6(x-1) \rightarrow [\lim_{x \rightarrow 1} \sin^{-1} (x-1)/\]